an airplane flies eastward and always accelerates at a constant

An airplane flies eastward and always accelerates
Answer #1

given

u=28.1 m/s

v=49.1m/s

s=50100m

v^{2}-u^{2}=2as

a=\frac{v^{2}-u^{2}}{2s}

a=0.016m/s^{2}

in three significant figure

time of elasped

v-u=at

t=\frac{v-u}{a}

t=1312.5s=1310s

in three significant figure